# The Minimax Lower Bound of Kernel Stein Discrepancy Estimation

### v1

@@ Abstract
Kernel Stein discrepancies (KSDs) have emerged as a powerful tool for quantifying goodness-of-fit over the last decade, featuring numerous successful applications. To the best of our knowledge, all existing KSD estimators with known rate achieve $\sqrt{n}$-convergence. In this work, we present two complementary results (with different proof strategies), establishing that the minimax lower bound of KSD estimation is $n^{-1/2}$ and settling the optimality of these estimators. Our first result focuses on KSD estimation on $\mathbb{R}^{d}$ with the Langevin-Stein operator; our explicit constant for the Gaussian kernel indicates that the difficulty of KSD estimation may increase exponentially with the dimensionality $d$. Our second result settles the minimax lower bound for KSD estimation on general domains.
@@ 1 INTRODUCTION
Similarly, KSD uses the mean embeddings of the target and the sampling distribution, where the underlying kernel is chosen such that the mean embedding of the target distribution vanishes. On Euclidean spaces, one attractive property of the classical Langevin-Stein KSD is that the resulting GoF measure is agnostic of the normalization constant of the sampling distribution, which can be challenging to compute in applications. This independence has led to its widespread use and its extension to other domains. Applications include model validation (Gorham and Mackey, 2017; Futami et al., 2019; Hodgkinson et al., 2021; Wang et al., 2023), learning variational models (Liu and Wang, 2016, 2018; Chen et al., 2018, 2019; Korba et al., 2020, 2021), testing (Liu et al., 2016; Chwialkowski et al., 2016; Schrab et al., 2022; Baum et al., 2023; Hagrass et al., 2025), model comparison (Lim et al., 2019; Kanagawa et al., 2020), and model explainability (Sarvmaili et al., 2025). KSD has also successfully been applied on discrete spaces (Yang et al., 2018), Riemannian manifolds (Xu and Matsuda, 2020, 2021; Barp et al., 2022), Hilbert spaces (Wynne et al., 2025), point-processes (Yang et al., 2019), and graph data (Xu and Reinert, 2021).
@@
    We establish the minimax lower bound $n^{-1/2}$ of KSD estimation on $\mathbb{R}^{d}$ with continuous bounded translation-invariant characteristic kernels, with explicit constants for Gaussian kernels.
@@
    Following a different proof strategy, we obtain the same lower bound for KSD estimation on general domains.
@@ 2 NOTATIONS
In this section, we introduce our notations: $\mathbb{N}_{0}$, $\mathbb{N}_{>0}$, $\mathbb{R}$, $[n]$, $\{\{\cdot\}\}$, $(\cdot)^{\top}$, $\langle\cdot,\cdot\rangle_{2}$, $\|\cdot\|_{2}$, $\|\cdot\|_{\infty}$, $\mathbf{1}_{d}$, $\mathbf{x}<\mathbf{y}$, $\mathbf{A}^{-}$, $\nabla$, $\operatorname{supp}$, $\bar{S}$, $\overline{z}$, $\overline{\mathbf{z}}$, $\text{Re}(\cdot)$, $\text{Im}(\cdot)$, $\mathbf{A}^{*}$, $\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{C}^{d}}$, $\|\mathbf{x}\|_{\mathbb{C}^{d}}$, $|\mathbf{x}|$, $\{\mathbf{e}_{j}\}_{j=1}^{d}$, $\frac{\partial f}{\partial x_{j}}$, $\mathbf{x}^{\bm{\alpha}}$, $D^{\bm{\alpha}}f$, $B(\mathscr{H})$, $\operatorname{Span}$, $\mathcal{C}^{s}(\mathbb{R}^{d})$, $\mathcal{C}(\mathcal{X})$, $\mathcal{C}_{b}(\mathcal{X})$, $\mathcal{F}_{c}(\mathcal{X})$, $\mathcal{M}_{1}^{+}(\mathcal{X})$, $\mathcal{B}(\mathcal{X})$, $\lambda_{d}$, $\delta_{x}$, $\mathbb{E}_{P}[X]$, $P^{n}$, $\ll$, $\frac{\mathrm{d}Q}{\mathrm{d}P}$, $\operatorname{KL}(Q\|P)$, $M_{\bm{\alpha}}^{P}$, $\mathcal{N}(\bm{\mu},\bm{\Sigma})$, $\psi_{P}$, $\mathscr{H}_{k}$, $\mathscr{H}_{k}^{d}$, $\mathcal{O}$, $\Omega$, $\Theta$, $\mathcal{O}_{P}$.
@@
The set of natural numbers is written as $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$; the set of positive integers is denoted by $\mathbb{N}_{>0}$; $\mathbb{R}$ stands for reals. Let $n]\coloneq\{1,\ldots,n\}$ with $n\in\mathbb{N}_{>0}$. We write $\{\{\cdot\}\}$ for a multiset. The transpose of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is written as $\mathbf{v}^{\top}\in\mathbb{R}^{1\times d}$. The inner product of vectors $\mathbf{u}=(u_{j})_{j=1}^{d},\mathbf{v}=(v_{j})_{j=1}^{d}\in\mathbb{R}^{d}$ is $\langle\mathbf{u},\mathbf{v}\rangle_{2}=\sum_{j=1}^{d}u_{j}v_{j}$. The Euclidean norm of $\mathbf{x}\in\mathbb{R}^{d}$ is $\|\mathbf{x}\|_{2}=\sqrt{\langle\mathbf{x},\mathbf{x}\rangle_{2}}$; its supremum norm is $\|\mathbf{x}\|_{\infty}=\max_{j\in[d]}|x_{j}|$. The $d$-dimensional vector of ones is denoted by $\mathbf{1}_{d}$. For vectors $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, $\mathbf{x}<\mathbf{y}$ means that $x_{j}<y_{j}$ for all $j\in[d]$. The (Moore-Penrose) pseudo-inverse of a matrix $\mathbf{A}\in\mathbb{R}^{d_{1}\times d_{2}}$ is $\mathbf{A}^{-}\in\mathbb{R}^{d_{2}\times d_{1}}$. For a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$, let $\nabla_{\mathbf{x}}f(\mathbf{x})=\left(\frac{\partial f}{\partial x_{j}}(\mathbf{x})\right)_{j=1}^{d}\in\mathbb{R}^{d}$ ($\mathbf{x}\in\mathbb{R}^{d}$). The support of a function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ is $\operatorname{supp}(\varphi)=\overline{\{\mathbf{x}\in\mathbb{R}^{d}:\varphi(\mathbf{x})\neq 0\}}$, where $\bar{S}$ stands for the closure of the set $S$. The conjugate of a complex number $z=a+ib\in\mathbb{C}$ is denoted by $\overline{z}=a-ib$ with $i=\sqrt{-1}$; its real part is $\text{Re}(z)=a$, its complex part is $\text{Im}(z)=b$. On a vector $\mathbf{z}=(z_{j})_{j=1}^{d}\in\mathbb{C}^{d}$, conjugation, real part and complex part act coordinate-wise: $\overline{\mathbf{z}}=(\overline{z_{j}})_{j=1}^{d}$, $\text{Re}(\mathbf{z})=(\text{Re}(z_{j}))_{j=1}^{d}$, $\text{Im}(\mathbf{z})=(\text{Im}(z_{j}))_{j=1}^{d}$. The adjoint of a matrix $\mathbf{A}\in\mathbb{C}^{d_{1}\times d_{2}}$ is written as $\mathbf{A}^{*}\in\mathbb{C}^{d_{2}\times d_{1}}$. The inner product of vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{d}$ is $\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{C}^{d}}=\mathbf{y}^{*}\mathbf{x}$; $\|\mathbf{x}\|_{\mathbb{C}^{d}}\coloneq\sqrt{\langle\mathbf{x},\mathbf{x}\rangle_{\mathbb{C}^{d}}}$ for $\mathbf{x}\in\mathbb{C}^{d}$. Let $\mathbf{x}=(x_{i})_{i=1}^{d}\in\mathbb{R}^{d}$; we write $|\mathbf{x}|\coloneq\sum_{j\in[d]}|x_{j}|$. Let $\{\mathbf{e}_{j}\}_{j=1}^{d}\subset\mathbb{R}^{d}$ be the canonical basis of $\mathbb{R}^{d}$. For $f:\mathbb{R}^{d}\rightarrow\mathbb{C}$, we define $\frac{\partial f}{\partial x_{j}}(\mathbf{x})=\lim\limits_{h\rightarrow 0}\frac{f(\mathbf{x}+h\mathbf{e}_{j})-f(\mathbf{x})}{h}=\lim\limits_{h\rightarrow 0}\frac{\text{Re}\left(f(\mathbf{x}+h\mathbf{e}_{j})\right)-\text{Re}\left(f(\mathbf{x})\right)}{h}+i\lim\limits_{h\rightarrow 0}\frac{\text{Im}\left(f(\mathbf{x}+h\mathbf{e}_{j})\right)-\text{Im}\left(f(\mathbf{x})\right)}{h}$ as the partial derivative of $f$ on $x_{j}$, and $\nabla_{\mathbf{x}}f(\mathbf{x})=\left(\frac{\partial f}{\partial x_{j}}(\mathbf{x})\right)_{j=1}^{d}\in\mathbb{C}^{d}$ as the gradient of $f$ ($\mathbf{x}\in\mathbb{R}^{d}$). Let $\bm{\alpha}=(\alpha_{j})_{j=1}^{d}\in\mathbb{N}_{0}^{d}$ and $\mathbf{x}\in\mathbb{R}^{d}$. We write $\mathbf{x}^{\bm{\alpha}}\coloneq\prod_{j=1}^{d}x_{j}^{\alpha_{j}}$ and $D^{\bm{\alpha}}f\coloneq\frac{\partial^{|\bm{\alpha}|}f}{\partial\mathbf{x}^{\bm{\alpha}}}=\frac{\partial^{|\bm{\alpha}|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}$. Let $\mathscr{H}$ be a Hilbert space; $B(\mathscr{H})=\{f\in\mathscr{H}\,:\,\left\|f\right\|_{\mathscr{H}}\leq 1\}$ denotes its unit ball centered at the origin. For a set $S$ in a vector space, $\operatorname{Span}(S)$ stands for its linear hull. For $s\in\mathbb{N}_{0}$, the space of $s$-times continuously differentiable real-valued functions on $\mathbb{R}^{d}$ is denoted by $\mathcal{C}^{s}(\mathbb{R}^{d})$. Let $\mathcal{X}$ be a topological space. The set of real-valued continuous functions on $\mathcal{X}$ is denoted by $\mathcal{C}(\mathcal{X})$. The subspace of $\mathcal{C}(\mathcal{X})$ consisting of bounded functions is denoted by $\mathcal{C}_{b}(\mathcal{X})$. The set of real-valued constant functions on $\mathcal{X}$ is denoted by $\mathcal{F}_{c}(\mathcal{X})$. The set of Borel probability measures on $\mathcal{X}$ is denoted by $\mathcal{M}_{1}^{+}(\mathcal{X})$, with $\mathcal{B}(\mathcal{X})$ standing for the Borel sigma algebra on $\mathcal{X}$. Let $\lambda_{d}$ denote the Lebesgue measure on $\mathbb{R}^{d}$. The Dirac measure centered at $x\in\mathcal{X}$ is denoted by $\delta_{x}$. The expectation of a random variable $X\sim P\in\mathcal{M}_{1}^{+}(\mathcal{X})$ is $\mathbb{E}_{P}[X]\ =\ \int_{\mathcal{X}}x\mathrm{d}P(x)$. The $n$-fold product measure of $P$ is denoted by $P^{n}=\otimes_{j=1}^{n}P$. Let $Q,P\in\mathcal{M}_{1}^{+}(\mathcal{X})$, and let $Q$ absolutely continuous w.r.t. $P$ ($Q\ll P$, with the corresponding Radon-Nikodym derivative denoted by $\frac{\mathrm{d}Q}{\mathrm{d}P}$), their Kullback–Leibler divergence is defined as $\operatorname{KL}(Q\|P)=\int_{\mathcal{X}}\ln\!\left(\frac{\mathrm{d}Q(x)}{\mathrm{d}P(x)}\right)\mathrm{d}Q(x)$. Given a probability measure $P\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$, we denote its moment of order $\bm{\alpha}\in\mathbb{N}_{0}^{d}$ as $M_{\bm{\alpha}}^{P}=\int_{\mathbb{R}^{d}}\mathbf{x}^{\bm{\alpha}}\mathrm{d}P(\mathbf{x})$. Normal random variables with mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$ are denoted by $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. The function $\psi_{P}(\bm{\omega})=\mathbb{E}_{P}[e^{i\left\langle{X,\bm{\omega}}\right\rangle_{2}}]$ is known as the characteristic function of $P$. A Hilbert space of functions $f:\mathcal{X}\to\mathbb{R}$ is a reproducing kernel Hilbert space (RKHS) $\mathscr{H}_{k}$ associated to a kernel $k\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ if $k(\cdot,x)\in\mathscr{H}_{k}$ for all $x\in\mathcal{X}$ and the reproducing property $f(x)=\langle f,k(\cdot,x)\rangle_{\mathscr{H}_{k}}$ holds for all $f\in\mathscr{H}_{k}$ and all $x\in\mathcal{X}$.222k(,x)k($\cdot$,x) denotes the function $x^{\prime}\in\mathcal{X}\mapsto k(x^{\prime},x)\in\mathbb{R}$ while keeping $x\in\mathcal{X}$ fixed. Let $\mathscr{H}_{k}^{d}=\mathscr{H}_{k}\times\cdots\times\mathscr{H}_{k}$ be the product RKHS with inner product $\langle\mathbf{f},\mathbf{g}\rangle_{\mathscr{H}_{k}^{d}}=\sum_{j=1}^{d}\langle f_{j},g_{j}\rangle_{\mathscr{H}_{k}}$ for $\mathbf{f}=(f_{j})_{j=1}^{d},\mathbf{g}=(g_{j})_{j=1}^{d}\in\mathscr{H}_{k}^{d}$. For positive sequences $(a_{n})_{n=1}^{\infty}$ and $\left(b_{n}\right)_{n=1}^{\infty}$, (i) $a_{n}=\mathcal{O}(b_{n})$ if there exist $C>0$ and $n_{0}\in\mathbb{N}_{>0}$ such that $a_{n}\leq Cb_{n}$ for all $n\geq n_{0}$, (ii) $a_{n}=\Omega(b_{n})$ if $b_{n}=\mathcal{O}(a_{n})$, (iii) $a_{n}=\Theta(b_{n})$ if $a_{n}=\mathcal{O}(b_{n})$ and $b_{n}=\mathcal{O}(a_{n})$. For a sequence of independent identically distributed (i.i.d.) real-valued random variables $(X_{n})_{n=1}^{\infty}$, $X_{n}\sim P$ and a sequence of positive reals $(a_{n})_{n=1}^{\infty}$ ($a_{n}>0$ for all $n$), $X_{n}=\mathcal{O}_{P}(a_{n})$ means that $\left(\frac{X_{n}}{a_{n}}\right)_{n=1}^{\infty}$ is bounded in probability.

### v2

@@ Abstract
Kernel Stein discrepancies (KSDs) have emerged as a powerful tool for quantifying goodness-of-fit over the last decade, featuring numerous successful applications. To the best of our knowledge, all existing KSD estimators with known rate achieve $\sqrt{n}$-convergence. In this work, we present two complementary results (with different proof strategies), establishing that the minimax lower bound of KSD estimation is $n^{-1/2}$ and settling the optimality of these estimators. Our first result focuses on KSD estimation on $\mathbb{R}^{d}$ with the Langevin-Stein operator; our explicit constant for the Gaussian base kernel indicates that the difficulty of KSD estimation may increase exponentially with the dimensionality $d$. Our second result settles the minimax lower bound for KSD estimation on general domains.
@@ 1 INTRODUCTION
Similarly, KSD uses the mean embeddings of the target and the sampling distribution, where the underlying kernel is chosen such that the mean embedding of the target distribution vanishes. On Euclidean spaces, one attractive property of the classical Langevin-Stein KSD is that the resulting GoF measure is agnostic of the normalization constant of the sampling distribution, which can be challenging to compute in applications. This independence has led to its widespread use and its extension to other domains. Applications include model validation (Gorham and Mackey, 2017; Futami et al., 2019; Hodgkinson et al., 2021; Wang et al., 2023), learning variational models (Liu and Wang, 2016, 2018; Chen et al., 2018, 2019; Korba et al., 2020, 2021), testing (Liu et al., 2016; Chwialkowski et al., 2016; Schrab et al., 2022; Baum et al., 2023; Hagrass et al., 2025), model comparison (Lim et al., 2019; Kanagawa et al., 2020), distribution compression (Li et al., 2024), and model explainability (Sarvmaili et al., 2025). KSD has also successfully been applied on discrete spaces (Yang et al., 2018), Riemannian manifolds (Xu and Matsuda, 2020, 2021; Barp et al., 2022), Hilbert spaces (Wynne et al., 2025), point-processes (Yang et al., 2019), and graph data (Xu and Reinert, 2021).
@@
    We establish the minimax lower bound $n^{-1/2}$ of KSD estimation on $\mathbb{R}^{d}$ with continuous bounded translation-invariant characteristic base kernels, with explicit constants for Gaussian base kernels.
@@
    Following a different proof strategy—by employing local perturbations—, we obtain the same lower bound for KSD estimation on general domains. Our imposed integrability conditions can be seen as a relaxation of the usual boundedness assumptions.
@@ 2 NOTATIONS
In this section, we introduce our notations: $\mathbb{N}_{0}$, $\mathbb{N}_{>0}$, $\mathbb{R}$, $[n]$, $\{\{\cdot\}\}$, $(\cdot)^{\top}$, $\langle\cdot,\cdot\rangle_{2}$, $\|\cdot\|_{2}$, $\|\cdot\|_{\infty}$, $\mathbf{1}_{d}$, $\mathbf{x}<\mathbf{y}$, $\mathbf{A}^{-}$, $\nabla$, $\operatorname{supp}$, $\bar{S}$, $\overline{z}$, $\overline{\mathbf{z}}$, $\text{Re}(\cdot)$, $\text{Im}(\cdot)$, $\mathbf{A}^{*}$, $\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{C}^{d}}$, $\|\mathbf{x}\|_{\mathbb{C}^{d}}$, $|\mathbf{x}|$, $\{\mathbf{e}_{j}\}_{j=1}^{d}$, $\frac{\partial f}{\partial x_{j}}$, $\mathbf{x}^{\bm{\alpha}}$, $D^{\bm{\alpha}}f$, $B(\mathscr{H})$, $\operatorname{Span}$, $\mathcal{C}^{s}(\mathbb{R}^{d})$, $\mathcal{C}(\mathcal{X})$, $\mathcal{C}_{b}(\mathcal{X})$, $\mathcal{M}_{1}^{+}(\mathcal{X})$, $\mathcal{B}(\mathcal{X})$, $\lambda_{d}$, $\delta_{x}$, $\mathbb{E}_{P}[X]$, $P^{n}$, $\ll$, $\frac{\mathrm{d}Q}{\mathrm{d}P}$, $\operatorname{KL}(Q\|P)$, $M_{\bm{\alpha}}^{P}$, $\mathcal{N}(\bm{\mu},\bm{\Sigma})$, $\psi_{P}$, $\mathscr{H}_{k}$, $\mathscr{H}_{k}^{d}$, $\mathcal{O}$, $\Omega$, $\Theta$, $\mathcal{O}_{P}$.
@@
The set of natural numbers is written as $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$; the set of positive integers is denoted by $\mathbb{N}_{>0}$; $\mathbb{R}$ stands for reals. Let $n]\coloneq\{1,\ldots,n\}$ with $n\in\mathbb{N}_{>0}$. We write $\{\{\cdot\}\}$ for a multiset. The transpose of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is written as $\mathbf{v}^{\top}\in\mathbb{R}^{1\times d}$. The inner product of vectors $\mathbf{u}=(u_{j})_{j=1}^{d},\mathbf{v}=(v_{j})_{j=1}^{d}\in\mathbb{R}^{d}$ is $\langle\mathbf{u},\mathbf{v}\rangle_{2}=\sum_{j=1}^{d}u_{j}v_{j}$. The Euclidean norm of $\mathbf{x}\in\mathbb{R}^{d}$ is $\|\mathbf{x}\|_{2}=\sqrt{\langle\mathbf{x},\mathbf{x}\rangle_{2}}$; its supremum norm is $\|\mathbf{x}\|_{\infty}=\max_{j\in[d]}|x_{j}|$. The $d$-dimensional vector of ones is denoted by $\mathbf{1}_{d}$. For vectors $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, $\mathbf{x}<\mathbf{y}$ means that $x_{j}<y_{j}$ for all $j\in[d]$. The (Moore-Penrose) pseudo-inverse of a matrix $\mathbf{A}\in\mathbb{R}^{d_{1}\times d_{2}}$ is $\mathbf{A}^{-}\in\mathbb{R}^{d_{2}\times d_{1}}$. For a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$, let $\nabla_{\mathbf{x}}f(\mathbf{x})=\left(\frac{\partial f}{\partial x_{j}}(\mathbf{x})\right)_{j=1}^{d}\in\mathbb{R}^{d}$ ($\mathbf{x}\in\mathbb{R}^{d}$). The support of a function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ is $\operatorname{supp}(\varphi)=\overline{\{\mathbf{x}\in\mathbb{R}^{d}:\varphi(\mathbf{x})\neq 0\}}$, where $\bar{S}$ stands for the closure of the set $S$. The conjugate of a complex number $z=a+ib\in\mathbb{C}$ is denoted by $\overline{z}=a-ib$ with $i=\sqrt{-1}$; its real part is $\text{Re}(z)=a$, its complex part is $\text{Im}(z)=b$. On a vector $\mathbf{z}=(z_{j})_{j=1}^{d}\in\mathbb{C}^{d}$, conjugation, real part and complex part act coordinate-wise: $\overline{\mathbf{z}}=(\overline{z_{j}})_{j=1}^{d}$, $\text{Re}(\mathbf{z})=(\text{Re}(z_{j}))_{j=1}^{d}$, $\text{Im}(\mathbf{z})=(\text{Im}(z_{j}))_{j=1}^{d}$. The adjoint of a matrix $\mathbf{A}\in\mathbb{C}^{d_{1}\times d_{2}}$ is written as $\mathbf{A}^{*}\in\mathbb{C}^{d_{2}\times d_{1}}$. The inner product of vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{d}$ is $\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{C}^{d}}=\mathbf{y}^{*}\mathbf{x}$; $\|\mathbf{x}\|_{\mathbb{C}^{d}}\coloneq\sqrt{\langle\mathbf{x},\mathbf{x}\rangle_{\mathbb{C}^{d}}}$ for $\mathbf{x}\in\mathbb{C}^{d}$. Let $\mathbf{x}=(x_{i})_{i=1}^{d}\in\mathbb{R}^{d}$; we write $|\mathbf{x}|\coloneq\sum_{j\in[d]}|x_{j}|$. Let $\{\mathbf{e}_{j}\}_{j=1}^{d}\subset\mathbb{R}^{d}$ be the canonical basis of $\mathbb{R}^{d}$. For $f:\mathbb{R}^{d}\rightarrow\mathbb{C}$, we define $\frac{\partial f}{\partial x_{j}}(\mathbf{x})=\lim\limits_{h\rightarrow 0}\frac{f(\mathbf{x}+h\mathbf{e}_{j})-f(\mathbf{x})}{h}=\lim\limits_{h\rightarrow 0}\frac{\text{Re}\left(f(\mathbf{x}+h\mathbf{e}_{j})\right)-\text{Re}\left(f(\mathbf{x})\right)}{h}+i\lim\limits_{h\rightarrow 0}\frac{\text{Im}\left(f(\mathbf{x}+h\mathbf{e}_{j})\right)-\text{Im}\left(f(\mathbf{x})\right)}{h}$ as the partial derivative of $f$ on $x_{j}$, and $\nabla_{\mathbf{x}}f(\mathbf{x})=\left(\frac{\partial f}{\partial x_{j}}(\mathbf{x})\right)_{j=1}^{d}\in\mathbb{C}^{d}$ as the gradient of $f$ ($\mathbf{x}\in\mathbb{R}^{d}$). Let $\bm{\alpha}=(\alpha_{j})_{j=1}^{d}\in\mathbb{N}_{0}^{d}$ and $\mathbf{x}\in\mathbb{R}^{d}$. We write $\mathbf{x}^{\bm{\alpha}}\coloneq\prod_{j=1}^{d}x_{j}^{\alpha_{j}}$ and $D^{\bm{\alpha}}f\coloneq\frac{\partial^{|\bm{\alpha}|}f}{\partial\mathbf{x}^{\bm{\alpha}}}=\frac{\partial^{|\bm{\alpha}|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}$. Let $\mathscr{H}$ be a Hilbert space; $B(\mathscr{H})=\{f\in\mathscr{H}\,:\,\left\|f\right\|_{\mathscr{H}}\leq 1\}$ denotes its unit ball centered at the origin. For a set $S$ in a vector space, $\operatorname{Span}(S)$ stands for its linear hull. For $s\in\mathbb{N}_{0}$, the space of $s$-times continuously differentiable real-valued functions on $\mathbb{R}^{d}$ is denoted by $\mathcal{C}^{s}(\mathbb{R}^{d})$. Let $\mathcal{X}$ be a topological space. The set of real-valued continuous functions on $\mathcal{X}$ is denoted by $\mathcal{C}(\mathcal{X})$. The subspace of $\mathcal{C}(\mathcal{X})$ consisting of bounded functions is denoted by $\mathcal{C}_{b}(\mathcal{X})$. The set of Borel probability measures on $\mathcal{X}$ is denoted by $\mathcal{M}_{1}^{+}(\mathcal{X})$, with $\mathcal{B}(\mathcal{X})$ standing for the Borel sigma algebra on $\mathcal{X}$. Let $\lambda_{d}$ denote the Lebesgue measure on $\mathbb{R}^{d}$. The Dirac measure centered at $x\in\mathcal{X}$ is denoted by $\delta_{x}$. The expectation of a random variable $X\sim P\in\mathcal{M}_{1}^{+}(\mathcal{X})$ is $\mathbb{E}_{P}[X]\ =\ \int_{\mathcal{X}}x\mathrm{d}P(x)$. The $n$-fold product measure of $P$ is denoted by $P^{n}=\otimes_{j=1}^{n}P$. Let $Q,P\in\mathcal{M}_{1}^{+}(\mathcal{X})$, and let $Q$ absolutely continuous w.r.t. $P$ ($Q\ll P$, with the corresponding Radon-Nikodym derivative denoted by $\frac{\mathrm{d}Q}{\mathrm{d}P}$), their Kullback–Leibler divergence is defined as $\operatorname{KL}(Q\|P)=\int_{\mathcal{X}}\ln\!\left(\frac{\mathrm{d}Q(x)}{\mathrm{d}P(x)}\right)\mathrm{d}Q(x)$. Given a probability measure $P\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$, we denote its moment of order $\bm{\alpha}\in\mathbb{N}_{0}^{d}$ as $M_{\bm{\alpha}}^{P}=\int_{\mathbb{R}^{d}}\mathbf{x}^{\bm{\alpha}}\mathrm{d}P(\mathbf{x})$. Normal random variables with mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$ are denoted by $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. The function $\psi_{P}(\bm{\omega})=\mathbb{E}_{P}[e^{i\left\langle{X,\bm{\omega}}\right\rangle_{2}}]$ is known as the characteristic function of $P$. A Hilbert space of functions $f:\mathcal{X}\to\mathbb{R}$ is a reproducing kernel Hilbert space (RKHS) $\mathscr{H}_{k}$ associated to a kernel $k\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ if $k(\cdot,x)\in\mathscr{H}_{k}$ for all $x\in\mathcal{X}$ and the reproducing property $f(x)=\langle f,k(\cdot,x)\rangle_{\mathscr{H}_{k}}$ holds for all $f\in\mathscr{H}_{k}$ and all $x\in\mathcal{X}$.222k(,x)k($\cdot$,x) denotes the function $x^{\prime}\in\mathcal{X}\mapsto k(x^{\prime},x)\in\mathbb{R}$ while keeping $x\in\mathcal{X}$ fixed. Let $\mathscr{H}_{k}^{d}=\mathscr{H}_{k}\times\cdots\times\mathscr{H}_{k}$ be the product RKHS with inner product $\langle\mathbf{f},\mathbf{g}\rangle_{\mathscr{H}_{k}^{d}}=\sum_{j=1}^{d}\langle f_{j},g_{j}\rangle_{\mathscr{H}_{k}}$ for $\mathbf{f}=(f_{j})_{j=1}^{d},\mathbf{g}=(g_{j})_{j=1}^{d}\in\mathscr{H}_{k}^{d}$. For positive sequences $(a_{n})_{n=1}^{\infty}$ and $\left(b_{n}\right)_{n=1}^{\infty}$, (i) $a_{n}=\mathcal{O}(b_{n})$ if there exist $C>0$ and $n_{0}\in\mathbb{N}_{>0}$ such that $a_{n}\leq Cb_{n}$ for all $n\geq n_{0}$, (ii) $a_{n}=\Omega(b_{n})$ if $b_{n}=\mathcal{O}(a_{n})$, (iii) $a_{n}=\Theta(b_{n})$ if $a_{n}=\mathcal{O}(b_{n})$ and $b_{n}=\mathcal{O}(a_{n})$. For a sequence of independent identically distributed (i.i.d.) real-valued random variables $(X_{n})_{n=1}^{\infty}$, $X_{n}\sim P$ and a sequence of positive reals $(a_{n})_{n=1}^{\infty}$ ($a_{n}>0$ for all $n$), $X_{n}=\mathcal{O}_{P}(a_{n})$ means that $\left(\frac{X_{n}}{a_{n}}\right)_{n=1}^{\infty}$ is bounded in probability.